OpenStudy (anonymous):

which expression is a cube root of 2i a. cube root of 2 (cos (210°) + i sin (210°)) b. cube root of 2 (cos (260°) + i sin (260°)) c. cube root of 2 (cos (90°) + i sin (90°)) d. cube root of 2 (cos (60°) + i sin (60°))

OpenStudy (anonymous):

Do you know demoivres?

OpenStudy (anonymous):

No I dont

OpenStudy (anonymous):

Ok so cos (theta) + isin(theta)=\[e ^{thetai}\]

OpenStudy (anonymous):

so first we express i in terms of cos(theta) +isin(theta

OpenStudy (anonymous):

how? like 2cos(theta) +2 i sin (theta) ?

OpenStudy (anonymous):

Ok I got D

OpenStudy (anonymous):

Ill draw it

jimthompson5910 (jim_thompson5910):

D is not correct, but it's close

jimthompson5910 (jim_thompson5910):

oh wait, lol wrong mode, my bad

OpenStudy (anonymous):

lol nice 99 dude ^

jimthompson5910 (jim_thompson5910):

one sec, I gotta redo

OpenStudy (anonymous):

so its D?

OpenStudy (anonymous):

Yea do you want me to teach you Demoivres ?

OpenStudy (anonymous):

yeah please. How did you get that?

OpenStudy (anonymous):

Awesome so i on the imaginary axis is (1, pi/2)

OpenStudy (anonymous):

so we can write i in terms of polar coordinates

OpenStudy (anonymous):

as cos (pi/2) +isin(pi/2)

OpenStudy (anonymous):

oh okay I think I got it now . Thanks

jimthompson5910 (jim_thompson5910):

The problem is this if z = cube root of 2 (cos (60°) + i sin (60°)) then z = cube root of 2 (cos (60°) + i sin (60°)) z^3 = (cube root of 2)^3 * ( cos(60*3) + i*sin(60*3) ) z^3 =2 * ( cos(180) + i*sin(180) ) but that will land you on the complex number 2+0i which is NOT 0+2i

jimthompson5910 (jim_thompson5910):

so this shows us that cube root of 2 (cos (60°) + i sin (60°)) is not a cube root of 2 (cos (90°) + i sin (90°))

jimthompson5910 (jim_thompson5910):

oops I meant to say -2+0i, but still, that's not 0+2i

OpenStudy (anonymous):

??? why did you multiply by 3?

jimthompson5910 (jim_thompson5910):

because if z = r( cos(theta) + i*sin(theta) ), then z^n = r^n*( cos(n*theta) + i*sin(n*theta) ) by De Moivre's theorem

jimthompson5910 (jim_thompson5910):

in this case, n = 3 because we're looking at cube roots

OpenStudy (anonymous):

ummm no

jimthompson5910 (jim_thompson5910):

yes, z is a cube root of z^3

jimthompson5910 (jim_thompson5910):

I'm checking answer choice D by cubing it You should be able to cube choice D and get 2 (cos (90°) + i sin (90°))